OpenStudy (anonymous):

(-1, 3) and perpendicular to y = - 1/3x + 7.

OpenStudy (anonymous):

then what ??

OpenStudy (anonymous):

y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

thats the answers i dont know how to do it

hero (hero):

To solve this, all you need to know is: 1. The product of the slopes of two perpendicular lines is -1. This means that if the slope of the first line is -1/3, then the slope of the second line must be 3 because -1/3 times 3 = -1. 2. y = mx + b; This is the general formula for a line in slope-intercept form. You need this to find the equation of the second line. You were given the point (-1,3) included on that line and we figured out that the slope of the second line is 3. 3. Plug in the given point and the slope of the second line, then solve for b: 3 = 3(-1) + b 3 = -3 + b 3 + 3 = b 6 = b 4. Now write the equation of the second line in point slope form. When you write it in point slope form, you only include the values of m and b: y = 3x + 6